A large law firm uses an average of 51 boxes of copier paper per day. The firm operates 260 days per year. Storage and handling costs for the paper are $21 per year per box, and it costs $49 to order and receive a shipment of paper.

On average, how many boxes of paper are in inventory? calculate to 4 decimal places
Answer #1

Given are following data :

Annual demand of copier paper = D = 51 / day x 260 days = 13260

Ordering cost = Co = $49

Annual unit storage and handling cost = Ch = $21

Therefore,

Economic Order Quantity ( EOQ )

= Square root ( 2 x Co x D/ Ch )

= Square root ( 2 x 49 x 13260 / 21 )

= 248.75

Average number of boxes of paper in inventory = EOQ / 2 = 248.75/2 = 124.3750

AVERAGE NUMBE ROF BOXES OF PAPER IN INVENTORY = 124.3750